Solve the following linear system for x using Cramer's rule show work
x+2y-3z=-22
2s-6y+8z=74
-x-2y+4z=29

Answer by Alan3354(69443)

x+2y-3z=-22
2s-6y+8z=74
-x-2y+4z=29
-------------
Find the determinant first.
1 2 -3
2 -6 8
-1 -2 4
-----------
1*(-24 + 16) -2*(8 + 8) -3*(-4 -6)
= -8 -32 +30
= -10
-------------
x =
2 -3 -22
-6 8 74
-2 4 29
divided by the Det
-10x = 2*(232 - 296) +3*(-174 + 148) -22*(-24 + 16)
-10x = -128 -78 +176
-10x = -30
x = 3
-------
10y =
1 -3 -22
2 8 74
-1 4 29
-----------
10y = 1*(232-296) + 3*(58+74) -22*(8+8)
10y = -64 + 396 -352
10y = -20
y = -2
-------------
-10z =
1 2 -22
2 -6 74
-1 -2 29
-----------
-10z = 1*(-174 + 148) - 2*(58 + 74) -22*(-4 -6)
-10z = -26 -264 +220
-10z = -70
z = 7
I'm working on an Excel program to do this.